# Efficient Transformer Encoders for Mask2Former-style models

### v1

@@ Abstract
Vision transformer based models bring significant improvements for image segmentation tasks. Although these architectures offer powerful capabilities irrespective of specific segmentation tasks, their use of computational resources can be taxing on deployed devices. One way to overcome this challenge is by adapting the computation level to the specific needs of the input image rather than the current one-size-fits-all approach. To this end, we introduce ECO-M2F or EffiCient TransfOrmer Encoders for Mask2Former-style models. Noting that the encoder module of M2F-style models incur high resource-intensive computations, ECO-M2F provides a strategy to self-select the number of hidden layers in the encoder, conditioned on the input image. To enable this self-selection ability for providing a balance between performance and computational efficiency, we present a three step recipe. The first step is to train the parent architecture to enable early exiting from the encoder. The second step is to create an derived dataset of the ideal number of encoder layers required for each training example. The third step is to use the aforementioned derived dataset to train a gating network that predicts the number of encoder layers to be used, conditioned on input image. Additionally, to change the computational-accuracy trade-off, only steps two and three need to be repeated which significantly reduces retraining time. Experiments on the public datasets show that the proposed approach reduces expected encoder computational cost while maintaining performance, adapts to various user compute resources, is flexible in architecture configurations, and can be extended beyond the segmentation task to object detection.
@@ 1 Introduction
Figure 1: Comparison to prior works. Instead of conventional M2F-style architecture that provides “one-size-fits-all” solution, our method ECO-M2F focuses on training such models in order to directly run at various resource encoder depths by leveraging a gating function. Here, B, E, D, and G denote the backbone, encoder, decoder, and (our proposed) gating network, respectively.
Figure 2: Histogram of images achieving best panoptic segmentation by number of encoder layers.
@@
With the advent of powerful universal image segmentation architectures [6, 5, 16, 11], it is highly desirable to prioritize the computational efficiency of these architectures for their enhanced scalability, e.g., use on resource-limited edge devices. These architectures are extremely useful in tackling instance [12], semantic [29], and panoptic [19] segmentation tasks using one generalized architecture, owing to the transformer-based [31] modules. These universal architectures leverage DEtection TRansformers or DETR-style [2] modules and represent both stuff and things categories [19] using general feature tokens. This is an incredible advantage over preceding segmentation methods [26, 15, 40] in literature that require careful considerations in design specifications. Hence, these segmentation architectures reduce the need for task-specific choices which favor performance of one task over the other [5].
@@
State-of-the-art models for universal segmentation like Mask2former (M2F) [5] are built on the key idea inspired from DETR: “mask” classification is versatile enough to address both semantic- and instance-level segmentation tasks. However, the problem of efficient M2F-style architectures have been under-explored. With backbone architectures (e.g., Resnet-50 [13], SWIN-Tiny [22]), [20] showed that DETR-style models incur the highest computations from the transformer encoder due to maintaining full length token representations from multi-scale backbone features. While existing works like [20, 24] primarily focus on scaling the input token to improve efficiency, this approach often neglects other aspects of model optimization and leads to a “one-size-fits-all” solution (Figure 2). This limitation leaves significant room for further efficiency improvements.
@@
Given this growing importance of M2F-style architectures and indispensable need for efficiency for real-world deployment, we introduce ECO-M2F or ‘EffiCient TransfOrmer Encoders’ for M2F-style architectures. Our key idea comes from our observation made on the training set of COCO [21] and Cityscapes [7] dataset demonstrated in Fig. 2. We plot a histogram of the number of hidden encoder layers that produces the best panoptic segmentation quality [19] for each image. It can be seen that not all images require the use of all $K$ hidden layers of the transformer encoder in order to achieve the maximum panoptic segmentation quality [19]. With this insight, we propose to create a dynamic transformer encoder that economically uses the hidden layers, guided by a gating network that can select different depths for different images.
@@
To achieve the aforementioned ability, ECO-M2F leverages the well-studied early exiting strategy [28, 37, 23, 33, 17, 41, 30, 27, 44] to create stochastic depths for the transformer encoder to improve inference efficiency. Previous exit mechanisms have primarily relied on confidence scores or uncertainty scores, typically applied in classification tasks. However, implementing such mechanisms in our context would necessitate the inclusion of a decoder and a prediction head to generate a reliable confidence score. This additional complexity introduces a significant number of FLOPs, rendering it impractical for our purposes. By contrast, ECO-M2F provides a three-step training recipe that can be used to customize the transformer encoder on the fly given the input image. Step A involves training the parent model to be dynamic by allowing stochastic depths at the transformer encoder. Using the fact that the transformer encoder maintains the token length of the input throughout the hidden layers constant, Step B involves creating a Derived dataset from the training dataset whose each sample contains a pair of image and layer number that provides the highest segmentation quality. Finally, Step C involves training a Gating Network using the derived dataset, whose function is to decide the number of layers to be used given the input image.
The key contributions of ECO-M2F are multifold. First, given a trained M2F-style architecture, ECO-M2F fine-tunes the model into one that allows the ability to randomly exit from the encoder by leveraging the fact that the token length remains constant in the hidden layers. Second, it introduces an accessory to the parent architecture via the Gating network that provides the ability to smartly use the encoder layers. Using this module, ECO-M2F enables the parent architecture to decide the optimal amount of layers for the given input without any performance degradation as well as any confidence threshold (unlike prior early exiting strategies). Third, as a result of our Gating network module’s training strategy, ECO-M2F can adapt the parent architecture to varying computational budgets using only Step C. On COCO [21] dataset, the computational cost of Step B and C are only $\sim$6and $\sim$2.5of Step A cost, respectively. Finally, ECO-M2F can also incorporate recent advances [20] in making transformer encoder efficient using token length scaling, bringing best of the both methods in pushing the limits of the efficiency. To summarize, we make the following contributions:
  • We present a dynamic transformer encoder ECO-M2F for M2F-style universal segmentation that maintains performance but reduces the computational load.
  • ECO-M2F consists of a novel training recipe that leverages input image based early exiting (in Step A), creating a derived dataset (based on training set segmentation performance in Step B), which in turn is used to train a gating function (using Step C) that allows adapting the number of hidden layers to reduce computations.
  • Extensive experiments show that ECO-M2F improves the overall performance-efficiency trade-off, and adaptable to diverse architecture settings and can be extended beyond segmentation to the detection task.
@@ 2 Related Works
With the rise of transformers [31], researchers are increasingly interested in creating image segmentation models that work effectively in various settings, without requiring segmentation type specific modifications to the model itself. Building on DETR [2], multiple universal segmentation architectures were proposed [6, 5, 16, 11] that use transformer decoder to predict masks for each entity in the input image. However, despite the significant progress in overall performance across various tasks, these models still face challenges in deployment on resource-constrained devices. Current emphasis [4, 10, 14, 15, 1, 38, 43, 42] for efficiency for image segmentation has mostly been on specialized architectures tailored to a single segmentation task. Unlike these preceding works, ECO-M2F makes no such assumption on the segmentation task and addresses the limitation of inefficiency in M2F-style universal architectures that are task-agnostic.
@@
Recent works on early exiting [32, 39, 28, 37, 23, 33, 17, 41, 30, 27, 44] aim to boost inference efficiency for large transformers. Some works [37, 23, 28] used early exiting for classification tasks along with manually chosen confidence threshold in vision transformers. For example, [37] proposed an early exiting framework for classification task ViTs combining heterogeneous task heads. Similarly, [28] proposed an early exiting strategy for vision-language models by measuring layer-wise similarities by checking multiple times to exit early. Applying early exiting solely to the encoder (like [37]) is infeasible due to the dependency on separate decoders, leading to an unacceptable optimization load. In contrast, methods like [28] suffer from redundant computations for exit decisions at all possible choices, hindering efficient resource allocation. In contrast, ECO-M2F only trains one decoder for all possible exit routes, as well as uses a gating module to decide the number of encoder layers required for the model depending on the input image.
@@ 3 Proposed Methodology: ECO-M2F
Model preliminaries.
We first review the meta-architecture of M2F [5] upon which ECO-M2F is based, along with the notation. This class of models contains
@@
For brevity, we collectively refer to the operations in the transformer decoder and segmentation head together as $h(\cdot)$. Thus, the output of the meta-architecture with $K$ encoder layers (a predicted mask $\tilde{\bm{y}}_{K}^{(i)}$ and corresponding label ~$\tilde{\ell}_{K}^{(i)}$) can be written as
@@
Here, the operation $\circ$ represents function composition, e.g., $g\circ f(x)=g(f(x))$ and subscript denotes output predicted using $K$ encoder layers. With $\{\bm{y}^{(i)},\ell^{(i)}\}$ as the pair of ground truth segmentation map and corresponding label of image ${\bm{x}}^{(i)}$, the final loss [5] is computed as
@@
    Model suitability for early exiting. Traditional early exiting techniques [28, 37, 23, 33, 17, 41, 30, 27, 44] often face challenges in maintaining satisfactory performance levels at potential exit points throughout the neural network. We recognize the importance of a model architecture that not only allows for early exiting but also ensures that the performance remains consistently high. Therefore, we aim to develop a model that not only permits early exits but also for which the accuracy steadily improves as the network delves deeper into its architecture. By prioritizing this aspect, we seek to establish a framework where early exiting does not compromise the overall performance of the model.
@@
    Efficient and effective gating network for optimal exit decision making. The efficacy of an early exiting strategy heavily depends on the ability to make informed exit decisions. A gating network must strike a delicate balance, minimizing computational overhead while effectively identifying components that can be bypassed without compromising accuracy. Our objective is to design a lightweight yet powerful gating mechanism capable of discerning optimal exit points within the model architecture.
@@
    Dynamic control mechanism for cost-performance trade-off. We require a mechanism with the ability to adaptively regulate the balance between computational cost and performance according to user-defined priorities. Such a mechanism empowers the model to exit at the optimal layer based on specific needs and desired outcomes, ensuring efficient resource allocation and maximizing utility in various application scenarios, particularly in resource-constrained environments like edge computing or real-time applications.
Figure 3: ECO-M2F framework. During the model pre-processing phase, we train the model to exit stochastically at $K$ potential exits using Step A. This is followed by Step B, where we use this model to perform inference on the training images at each exit to create a dataset $\mathcal{D}$. In the model adaptation phase, we perform Step C to establish a gating target based on the computational budget and train a lightweight gating network. During inference, the network adheres to the gating network’s output and exits at its designated output layer.
Training and Inference overview.
As shown in Fig. 3, the training phase of ECO-M2F involves following three main steps:
  1. $\bullet$
    Step A: Train parent model for early exit via the transformer encoder.
  2. $\bullet$
  3. $\bullet$
@@
3.1 Step A: Training the Model with Weighted Stochastic Depth
In this step, we enable the model to allow exiting at the encoder. To maintain consistently high performance at each exit point, we input each stochastic depth’s output to a shared transformer decoder. We then apply Eq. 2 to compute the loss $\mathcal{L}_{k}$ for each exit point $k$. However, we observe that direct training in this fashion does not encourage the model to use fewer layers to extract and prioritize informative representations, as shown in Tab. 6. To address this, we introduce a set of coefficients $\alpha_{k}$ to emphasize the quality of representations at later layers more, enabling earlier layers to also concentrate on producing effective intermediate representations. As the layer depth increases, the corresponding coefficient $\alpha_{k}$ grows, ensuring a progressively stricter standard for feature quality. The new loss function is then expressed as
@@ 3.1 Step A: Training the Model with Weighted Stochastic Depth
Figure 4: Intuition for Eq. 4. This figure shows that prioritizing PQ would need more encoder layers, and conversely, prioritizing lesser layers would result in poorer PQ. (Backbone: SWIN-T; training set).
3.2 Step B: Deriving the Gating Network Training Dataset
@@
To this end, we record the performance of the pre-trained stochastic depth model (obtained from Step A) at all potential exit points for each image within the training dataset and create a Derived dataset $\mathcal{D}$. Specifically, we associate the $i$-th input image ${\bm{x}}^{(i)}$ with a vector ${\bm{q}}^{(i)}$ of length $K$. Each element ${q}_{k}^{(i)}$ of ${\bm{q}}^{(i)}$ represents the predicted panoptic quality [19] upon exiting at the encoder layer $k$. Hence, each sample of $\mathcal{D}$ can be represented as $({\bm{x}}^{(i)},{\bm{q}}^{(i)})\in\mathcal{D}$.
3.3 Step C: Training for Gating Network
In this step, we train the gating network on dataset $\mathcal{D}$ (obtained from Step B) to self-select the number of encoder layers based on the input image. Ideally, this module should allow exiting at the encoder layer which would result in the highest quality segmentation map. With this in mind, we first establish the target exit for the gating network. Note that the panoptic quality generally increases with increasing encoder layers (see Fig. 4). However, we would like the gating network to prioritize increasing the panoptic quality while also reducing the number of layers (to reduce the overall computations). Consequently, we introduce a utility function expressed as the linear combination of segmentation quality and the depth of the network. This function is formulated as
@@ 3.3 Step C: Training for Gating Network
In consideration of having minimal impact on the computations due to the gating network, we use the output of the lowest resolution feature map $\bm{s}_{1}$ as input to the pooling operation. To optimize the gating network, we use the standard cross-entropy loss between the output logits ${\bm{g}}^{(i)}$ and the one-hot version of target exit $t^{(i)}$ as our training objective. During inference, the gating network identifies the layer with the highest predicted logits, i.e., $\operatorname*{arg\,max}_{k}(g_{k}^{(i)})$, as the optimal exit layer for image ${\bm{x}}^{(i)}$. Note that while there can be more complex choices for the gating network, our simple linear layer in Eq. 6 works well in our experiments.
Saving training costs through Step C.

### v2

# Image-Specific Adaptation of Transformer Encoders for Compute-Efficient Segmentation

@@ Abstract
Vision transformer-based models bring significant improvements to image segmentation tasks. Although these architectures offer powerful capabilities irrespective of specific segmentation tasks, their use of computational resources can be taxing on deployed devices. One way to overcome this challenge is by adapting the computation level to the specific needs of the input image rather than the current one-size-fits-all approach. To this end, we introduce ECO-M2F or EffiCient TransfOrmer Encoders for Mask2Former-style models. Noting that the encoder module of M2F-style models incur high resource-intensive computations, ECO-M2F provides a strategy to self-select the number of hidden layers in the encoder, conditioned on the input image. To enable this self-selection ability for providing a balance between performance and computational efficiency, we present a three-step recipe. The first step is to train the parent architecture to enable early exiting from the encoder. The second step is to create a derived dataset of the ideal number of encoder layers required for each training example. The third step is to use the aforementioned derived dataset to train a gating network that predicts the number of encoder layers to be used, conditioned on input images. Additionally, to change the computational-accuracy trade-off, only steps two and three need to be repeated which significantly reduces retraining time. Experiments on the public datasets show that the proposed approach reduces expected encoder computational cost while maintaining performance, adapts to various user compute resources, is flexible in architecture configurations, and can be extended beyond the segmentation task to object detection.
@@ 1 Introduction
With the advent of powerful universal image segmentation architectures [6, 5, 16, 11], it is highly desirable to prioritize the computational efficiency of these architectures for their enhanced scalability, e.g., use on resource-limited edge devices. These architectures are extremely useful in tackling instance [12], semantic [29], and panoptic [19] segmentation tasks using one generalized architecture, owing to the transformer-based [31] modules. These universal architectures leverage DEtection TRansformers or DETR-style [2] modules and represent both stuff and things categories [19] using general feature tokens. This is an incredible advantage over preceding segmentation methods [26, 15, 38] in the literature that require careful considerations in design specifications. Hence, these segmentation architectures reduce the need for task-specific choices that favor the performance of one task over the other [5].
Figure 1: Comparison to prior works. Instead of conventional M2F-style architecture that provides a “one-size-fits-all” solution, our method ECO-M2F trains models to run directlys at various resource encoder depths by leveraging a gating function. B, E, D, and G denote the backbone, encoder, decoder, and (our proposed) gating network, respectively.
Figure 2: (a) Histogram of images achieving best panoptic segmentation by the number of encoder layers. (b) Our method demonstrates superior performance and lower computational cost compared to the baseline models. (Dataset: Cityscapes; Backbone: SWIN-T)
@@
State-of-the-art models for universal segmentation like Mask2former (M2F) [5] are built on the key idea inspired by DETR: “mask” classification is versatile enough to address both semantic- and instance-level segmentation tasks. However, the problem of efficient M2F-style architectures has been underexplored. With backbone architectures (e.g., Resnet-50 [13], SWIN-Tiny [22]), [20] showed that DETR-style models incur the highest computations from the transformer encoder due to maintaining full-length token representations from multi-scale backbone features. While existing works like [20, 24] primarily focus on scaling the input token to improve efficiency, this approach often neglects other aspects of model optimization and leads to a “one-size-fits-all” solution (Fig. 1). This limitation leaves significant room for further efficiency improvements.
@@
Given this growing importance of M2F-style architectures and the indispensable need for efficiency for real-world deployment, we introduce ECO-M2F or ‘EffiCient TransfOrmer Encoders’ for M2F-style architectures. Our key idea comes from our observation made on the training set of COCO [21] and Cityscapes [7] dataset demonstrated in Fig. 2 (a). We plot a histogram of the number of hidden encoder layers that produces the best panoptic segmentation quality [19] for each image. It can be seen that not all images require the use of all $K$ hidden layers of the transformer encoder in order to achieve the maximum panoptic segmentation quality [19]. With this insight, we propose to create a dynamic transformer encoder that economically uses the hidden layers, guided by a gating network that can select different depths for different images.
@@
To achieve the aforementioned ability, ECO-M2F leverages the well-studied early exiting strategy [28, 35, 23, 33, 17, 39, 30, 27, 42] to create stochastic depths for the transformer encoder to improve inference efficiency. Previous exit mechanisms have primarily relied on confidence scores or uncertainty scores, typically applied in classification tasks. However, implementing such mechanisms in our context would necessitate the inclusion of a decoder and a prediction head to generate a reliable confidence score. This additional complexity introduces a significant number of FLOPs, rendering it impractical for our purposes. By contrast, ECO-M2F provides a three-step training recipe that can be used to customize the transformer encoder on the fly given the input image. Step A involves training the parent model to be dynamic by allowing stochastic depths at the transformer encoder. Using the fact that the transformer encoder maintains the token length of the input throughout the hidden layers constant, Step B involves creating a Derived dataset from the training dataset whose each sample contains a pair of images and layer number that provides the highest segmentation quality. Finally, Step C involves training a Gating Network using the derived dataset, whose function is to decide the number of layers to be used given the input image.
The key contributions of ECO-M2F are manifold:
  • Enhanced Encoder Flexibility: It fine-tunes pre-trained M2F-style architectures for early exits from the encoder, leveraging the constant token length in hidden layers.
  • Adaptive Gating Network Training: It introduces a Gating network to optimize encoder layer usage, allowing the architecture to determine the optimal layer count per input without performance loss or confidence thresholds.
  • Innovative Efficient Training: The Gating network’s training approach enables the architecture to adapt to varying computational budgets, with the cost of only Step C(about 2.5% of Step Aon COCO [21] dataset).
  • Integration and Expansion: ECO-M2F can incorporate recent advancements in token length scaling for transformer encoder efficiency and extend its application open-vocabulary segmentation and detection tasks.
  • Competitive Performance-Computational Trade-off: Our method strikes a competitive balance between performance and computational efficiency, and can even excel in both areas simultaneously, as illustrated in Fig. 2 (b).
@@ 2 Related Works
With the rise of transformers [31], researchers are increasingly interested in creating image segmentation models that work effectively in various settings, without requiring segmentation type-specific modifications to the model itself. Building on DETR [2], multiple universal segmentation architectures were proposed [6, 5, 16, 11] that use a transformer decoder to predict masks for each entity in the input image. However, despite the significant progress in overall performance across various tasks, these models still face challenges in deployment on resource-constrained devices. Current emphasis [4, 10, 14, 15, 1, 36, 41, 40] for efficiency for image segmentation has mostly been on specialized architectures tailored to a single segmentation task. Unlike these preceding works, ECO-M2Fmakes no such assumption on the segmentation task and addresses the limitation of inefficiency in M2F-style universal architectures that are task-agnostic.
@@
Recent works on early exiting [32, 37, 28, 35, 23, 33, 17, 39, 30, 27, 42] aim to boost inference efficiency for large transformers. Some works [35, 23, 28] used early exiting for classification tasks along with manually chosen confidence threshold in vision transformers. For example, [35] proposed an early exiting framework for classification task ViTs combining heterogeneous task heads. Similarly, [28] proposed an early exiting strategy for vision-language models by measuring layer-wise similarities by checking multiple times to exit early. Applying early exiting solely to the encoder (like [35]) is infeasible due to the dependency on separate decoders, leading to an unacceptable optimization load. In contrast, methods like [28] suffer from redundant computations for exit decisions at all possible choices, hindering efficient resource allocation. In contrast, ECO-M2Fonly trains one decoder for all possible exit routes, as well as uses a gating module to decide the number of encoder layers required for the model, depending on the input image.
@@ 3 Proposed Methodology: ECO-M2F
3.1 Model Preliminaries
We first review the meta-architecture of M2F [5] upon which ECO-M2F is based, along with the notation. This class of models contains:
@@ 3.1 Model Preliminaries
For brevity, we collectively refer to the operations in the transformer decoder and segmentation head together as $h(\cdot)$. Thus, the output of the meta-architecture with $K$ encoder layers (a predicted mask $\tilde{\bm{y}}_{K}^{(i)}$ and corresponding label $\tilde{\ell}_{K}^{(i)}$) can be written as
@@
Here, the operation $\circ$ represents function composition, e.g., $g\circ f(x)=g(f(x))$, and subscript denotes output predicted using $K$ encoder layers. With $\{\bm{y}^{(i)},\ell^{(i)}\}$ as the pair of ground truth segmentation map and corresponding label of image ${\bm{x}}^{(i)}$, the final loss [5] is computed as
@@ 3.2 Method Motivation
1. Model suitability for early exiting. Traditional early exiting techniques [28, 35, 23, 33, 17, 39, 30, 27, 42] often face challenges in maintaining satisfactory performance levels at potential exit points throughout the neural network. We recognize the importance of a model architecture that not only allows for early exiting but also ensures that the performance remains consistently high. Therefore, we aim to develop a model that not only permits early exits but also for which the accuracy steadily improves as the network delves deeper into its architecture. By prioritizing this aspect, we seek to establish a framework where early exiting does not compromise the overall performance of the model.
2. Efficient and effective gating network for optimal exit decision-making. The efficacy of an early exiting strategy heavily depends on the ability to make informed exit decisions. A gating network must strike a delicate balance, minimizing computational overhead while effectively identifying components that can be bypassed without compromising accuracy. Our objective is to design a lightweight yet powerful gating mechanism capable of discerning optimal exit points within the model architecture.
3. Dynamic control mechanism for cost-performance trade-off. We require a mechanism with the ability to adaptively regulate the balance between computational cost and performance according to user-defined priorities. Such a mechanism empowers the model to exit at the optimal layer based on specific needs and desired outcomes, ensuring efficient resource allocation and maximizing utility in various application scenarios, particularly in resource-constrained environments like edge computing or real-time applications.
Training and Inference Overview
Figure 3: ECO-M2F framework. During the model pre-processing phase, we train the model to exit stochastically at $K$ potential exits using Step A. Next, in Step B, we use this model to perform inference on the training images at each exit to create a dataset $\mathcal{D}$. In the model adaptation phase, we perform Step C to establish a gating target based on the computational budget and train a lightweight gating network. During inference, the network exists at the layer designated by the gating network.
As shown in Fig. 3, the training phase of ECO-M2F comprises three main steps:
@@
    Step A: Train the parent model for early exit via the transformer encoder.
@@
3.3 Training
Step A: Training with Weighted Stochastic Depth
In this step, we enable the model to allow exiting at the encoder. To maintain consistently high performance at each exit point, we input each stochastic depth’s output to a shared transformer decoder. We then apply Eq. 2 to compute the loss $\mathcal{L}_{k}$ for each exit point $k$. However, we observe that direct training in this fashion does not encourage the model to use fewer layers to extract and prioritize informative representations, as shown in the experiment results. To address this, we introduce a set of coefficients $\alpha_{k}$ to emphasize the quality of representations at later layers more, enabling earlier layers to also concentrate on producing effective intermediate representations. As the layer depth increases, the corresponding coefficient $\alpha_{k}$ grows, ensuring a progressively stricter standard for feature quality. The new loss function is then expressed as
@@ 3.3 Training
Step B: Deriving the Gating Network Training Set
@@
To this end, we record the performance of the pre-trained stochastic depth model (obtained from Step A) at all potential exit points for each image within the training dataset and create a Derived dataset $\mathcal{D}$. Specifically, we associate the $i$-th input image ${\bm{x}}^{(i)}$ with a vector ${\bm{q}}^{(i)}$ of length $K$. Each element ${q}_{k}^{(i)}$ of ${\bm{q}}^{(i)}$ represents the predicted panoptic quality [19] upon exiting at the encoder layer $k$. Hence, each sample of $\mathcal{D}$ can be represented as $({\bm{x}}^{(i)},{\bm{q}}^{(i)})$.
Step C: Training for Gating Network
In this step, we train the gating network on dataset $\mathcal{D}$ (obtained from Step B) to self-select the number of encoder layers based on the input image. Ideally, this module should allow exiting at the encoder layer, which would result in the highest quality segmentation map. With this in mind, we first establish the target exit for the gating network. Note that the panoptic quality generally increases with increasing encoder layers (see Fig. 4). However, we would like the gating network to prioritize increasing the panoptic quality while also reducing the number of layers (to reduce the overall computations). Consequently, we introduce a utility function expressed as a linear combination of segmentation quality and the depth of the network. This function is formulated as
@@
In consideration of having minimal impact on the computations due to the gating network, we use the output of the lowest resolution feature map $\bm{s}_{1}$ as input to the pooling operation. To optimize the gating network, we use the standard cross-entropy loss between the output logits ${\bm{g}}^{(i)}$ and the one-hot version of target exit $t^{(i)}$ as our training objective. During inference, the gating network identifies the layer with the highest predicted logits, i.e., $\operatorname*{arg\,max}_{k}(g_{k}^{(i)})$, as the optimal exit layer for image ${\bm{x}}^{(i)}$. Note that while there can be more complex choices for the gating network, our simple linear layer in Eq. 6 works well in experiments.
Figure 4: Intuition for Eq. 4. This figure shows that prioritizing PQ requires more encoder layers, while fewer layers lead to poorer PQ. (Backbone: SWIN-T; training set).
Saving Training Costs through Step C